Assuming that the skater has a speed of zero at position A, (Top of half pipe) which is 3.5m above the bottom of the half pipe.

Calculate the speed of the skater at position B (The middle bottom of the half pipe)
1
Expert's answer
2019-03-28T05:04:34-0400

Simply apply the law of conservation of energy: the potential energy transformed into kinetic energy:


"PE=KE,"

"mgh=\frac{1}{2}mv^2,"

"v=\sqrt{2gh}=\sqrt{2\cdot9.8\cdot3.5}=8.28\text{ m/s}."